# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ I.3 A glossary for aging dynamics
    Two relevant temperatures are $T_{\mathrm{g}}$ and $T_{\mathrm{m}}$. The glass transition occurs at $T_{\mathrm{g}}$ (which also is the critical temperature for the phase transition that separates the high-temperature paramagnetic phase from the spin-glass phase that becomes stable below $T_{\mathrm{g}}$). In a typical experimental setting, the system is quickly cooled from some very high temperature $T>T_{\mathrm{g}}$ to the measuring temperature $T_{\mathrm{m}}$.
@@
    Two relevant times are $t$ and $t_{\mathrm{w}}$. We set our initial time as the time for the temperature quench from $T>T_{\mathrm{g}}$ to $T_{\mathrm{m}}<T_{\mathrm{g}}$. The system is allowed to relax at $T_{\mathrm{m}}$ for the waiting time $t_{\mathrm{w}}$. Then, at time $t_{\mathrm{w}}$ the value of the externally applied magnetic field is varied (in the ZFC protocol the field is switched on at time $t_{\mathrm{w}}$; instead, in the TRM protocol the field is switched off). The magnetization is measured at a later time $t+t_{\mathrm{w}}$, giving us information about the sample’s reaction to the variation of the magnetic field at the earlier time $t_{\mathrm{w}}$.
@@
    We shall need two kind of averages. The thermal average of a quantity $\mathcal{A}$ will be denoted by $\langle\mathcal{A}\rangle$. The average over disorder, which is always performed after the thermal average, will be denoted by $\overline{\langle\mathcal{A}\rangle}$. Although the average over disorder is absolutely crucial for numerical simulations, experiments are usually carried out for just one sample. Due to self-averaging, this is usually a sound procedure under experimental conditions (see Sec. II.7).
@@
    The two-time correlation length $\zeta(t,t_{\mathrm{w}})$ describes the growth of order in the aging regime when the two times $t_{\mathrm{w}}$ and $t+t_{\mathrm{w}}$ are compared [157, 20, 44]. This quantity, defined in Sec. IV.2, will be vital for understanding temperature-varying protocols.
@@
    The experimentally accessible $\xi_{\mathrm{Zeeman}}$ introduced by Joh et al. [159] will be explained in Sec. 13. This length is related to the lowering of free-energy barriers because of the Zeeman effect and is non-linear in the magnetic field.
As explained in Sec. V, the relationship between the three length scales $\xi_{\mathrm{micro}}$, $\xi_{\mathrm{Zeeman}}$ and $\zeta$ is the key to the study of the non-linear response of a spin glass in a magnetic field. Although $\xi_{\mathrm{micro}}$ and $\xi_{\mathrm{Zeeman}}$ turn out to behave very similarly whenever they can be compared (which could suggest also referring to $\xi_{\mathrm{Zeeman}}$ as a coherence length), we shall follow the experimental tradition of referring to $\xi_{\mathrm{Zeeman}}$ as a correlation length.
@@ II.1 The Edwards-Anderson and Sherrington-Kirkpatrick models and overlap as an order parameter
By assuming that $\sum_{i,j}J_{ij}\epsilon_{ij}=0$ on all length scales of the system (where $J_{ij}$ is the strength of the quenched interaction between sites $i$ and $j$ and $\epsilon_{ij}$ is the probability of finding two interacting spins at those positions), Edwards and Anderson [104] introduced a very simplified model of classical spins (which will turn out to embed all relevant features of the problem), the Edwards-Anderson (EA) model. A different approach, based on the presence of potentially magnetized domains, had been proposed before by Adkins and Rivier [2]. The fact that the disordered interactions are distributed at random and cannot change in time is a crucial element of the model.
@@
where the primed sum runs over the pairs of nearest neighbors in a $D$-dimensional hypercubic lattice. The spins $\sigma_{i}$ can take different forms. For instance, in the Heisenberg spin glass, they are unit vectors in space. Unless otherwise specified, however, in this review, the EA model will always be considered with Ising spins that can only take the values ±$\pm 1$. The applied magnetic field $H$ will be considered to be zero, for now.
Notice that this Hamiltonian is written in natural units where $k_{\text{B}}=1$. To compare with experimental results, temperatures can be expressed in terms of $T_{\mathrm{g}}$ (more sophisticated scaling variables will be introduced later). As for the magnetic field, it can be estimated [33, 21], that $H=1$ roughly corresponds to $5\times 10^{4}$ Oe. Finally, the natural unit of time in simulations is the lattice sweep, which is equivalent to $\sim 1$ ps [218].
It is crucial that the quenched couplings, which do not change in the course of the dynamics, can have both signs. A possible choice is $J_{ij}=\pm 1$ with a probability of one half, or that they be normally distributed, with zero mean and unitary variance, $P\left(J_{ij}\right)\sim\exp(-{J_{ij}^{2}}/{2J^{2}})$. We will not discuss here the possibility of a non-zero expectation value for the couplings.
Because the quenched couplings can be both positive or negative and are distributed at random, some of the products of the couplings on a closed path of consecutive lattice links will be positive and some will be negative. This implies, in turn, that not all interactions can be satisfied at the same time; the system is frustrated, in the sense that all interactions can never be completely satisfied. The energy cannot be as low as it could be in a uniform system. Typically frustration leads to degeneracy: the low-energy states of such a system are not as good as those of a pure systems, but are numerous.
In the EA model, because the interactions between spins oscillate in sign as a function of separation, ferromagnetism does not appear at any relevant length scale. On the contrary, for $T$ lower than a critical $T_{\mathrm{g}}$, a spin-glass phase appears, with the spins staying close to fixed directions that appear to be random. This results in overall zero magnetization, but a non-zero square of the local magnetization and leads to the cusp in the susceptibility that was observed in experiments (see Sec. I.2).
@@
The absence of long-range magnetic order leads to the concept of spin overlap as an order parameter, which will be crucial for all further development of the field. One considers the spin $\sigma_{i}$ at site $i$ and defines:
$q_{i}\left(t,t+\tau\right)\equiv\lim_{\tau\to\infty}\left\langle\sigma_{i}\left(t\right)\sigma_{i}\left(t+\tau\right)\right\rangle,$ (4)
[…]
@@
This is the most common observable in numerical simulations. The concept of a non-zero overlap of configurations with zero magnetization is illustrated in Fig. 11.
@@
As with the overlap, we can associate a $P(q_{\text{l}})$ to this observable. The behavior of the overlap and the link overlap under spin flip is very different: consider two configurations of spins with the same disorder which differ in a domain of size $L$, where the spins have flipped, the link overlap provides an estimate for the surface of the domain, while the overlap will reflect its volume.
The EA spin glass, an Ising-like finite-dimensional model with quenched, frustrated interaction, cannot be solved exactly and cannot be dissected too deeply by analytical techniques. The experience of the Ising ferromagnet suggested introducing the mean-field, non-local Sherrington-Kirkpatrick (SK) model [260, 178].
@@
This mean-field model has a usual paramagnetic phase for high temperature. The delicate point, which will be the source of trouble and delight, is that, because the coupling are quenched, we have to compute the disorder average not of the partition function $Z$ (which would anneal the disorder degrees of freedom) but of its logarithm, $\ln Z$. That is, we average physical observables in a given disorder realization (a sample) and then we average these thermal expectation values over the disorder. To do this, EA and SK use the identity (the so-called replica trick)
@@
When the temperature decreases below a critical value $T_{\mathrm{g}}$ the correct solution found at high $T$ under the assumption of symmetry between replicas turns out to be undoubtedly wrong. Not only does it not agree with Monte Carlo simulations but, what is far more damning, it develops a negative entropy for low-enough $T$.
Kirkpatrick and Sherrington [178] refined their analysis two years after their first, groundbreaking paper [260]. They described the same situation, but with an analysis based on the assumption of replica symmetry (RS). Before this second SK paper, Thouless et al. [273] had introduced a different approach, avoiding the replica trick, that shed further light on the low-$T$ behavior of the SK model. Their so-called TAP equations, together with their many variations, have been for many years and continue to be important for the analysis of spin-glass physics.
@@
The above-described mean-field approximation is only correct above the so-called upper critical dimension $D_{\text{u}}$, which for spin glasses is $D_{\text{u}}=6$ [73]. On the other hand, below the lower critical dimension $D_{\text{l}}=2.5$ [124, 197], the low-temperature phase disappears. As a consequence, $D=3$ lies in the non-trivial region. There has been an intense quest for a successful extension of the Parisi theory to $D=3$ in the renormalization-group style [see for example 98, 95]. Because of the infinity of states that appear in Parisi’s solution for the mean-field theory, renormalizing the system is a daunting task. Today, after close to 50 years of studies, there still remains much need for clarification. The interested reader can start from De Dominicis et al. [98] and De Dominicis and Giardina [95], then continue with the following papers and included references: Bray and Moore [68, 69], Bray and Roberts [73], De Dominicis and Kondor [96, 97], Temesvári et al. [269], De Dominicis et al. [94], Temesvári [268].
@@ II.2 Main theoretical pictures for the low-temperature phase
  • Trivial-non-Trivial (TNT). In the droplet model (in the absence of a magnetic field) the probability distribution $P(q)$ of the overlap $q$ between two pure states, and the link overlap, are both trivial (i.e., $P(q)$ is composed of two Dirac delta functions and one Dirac delta function, respectively). In the TNT approach, $P(q)$ exhibits a non-trivial probability distribution, whereas the link overlap exhibits a trivial probability distribution [183, 234].
  • Replica Symmetry Breaking (RSB). In this scenario, which consists in extending the features of the mean-field solution to finite dimensions [203, 95], the spin glass has infinite pure states not related by any kind of symmetry. However, this infinite set of states can be organized in an ultrametric fashion. Furthermore, the excitations of the ground state are space-filling. Finally, in this theoretical approach the spin-glass phase is stable to small magnetic fields (i.e., there exists a phase transition between a paramagnetic phase and a spin-glass phase in the presence of a magnetic field).
Figure 12: The main quantitative difference between the competing theoretical pictures of spin-glass dynamics is related to the value of an exponent $\theta$, the replicon, see Eq. (23) and Marinari et al. [203, 202]. This exponent is non-zero for RSB and zero in the droplet picture, at least in the long-time limit. When represented as a function of a scaling variable $x$ [a properly rescaled $1/\xi(t_{\mathrm{w}},T)$, see 34], the values of $\theta$ obtained for several temperatures fall on a single curve. In addition we show a droplet extrapolation [dashed line, $\theta(x\!\to\!0)\!\propto\!x^{\iota}$, with $\iota\!\approx\!0.15$] and an RSB extrapolation [full line, $\theta(x\!\to\!0)\!>\!0$]. Even the droplet extrapolation, which has $\theta(x)\to 0$ as $x\to 0$, predicts sizable values of $\theta$ for the values of $x$ relevant to experimental length scales (black dots). Data taken from Baity-Jesi et al. [34].
@@ II.3 Phenomenological dynamical model on an ultrametric tree
There are no dynamics associated with pure states in that the free-energy barriers between them are infinite. Experimentally though, time-dependent phenomena are observed. It took a large amount of chutzpah to suggest that between the pure states there could exist states between which the free-energy barriers were finite and even more to suggest that these states between infinite barriers were self-similar: no matter the temperature, the same ultrametricity held for metastable states, which were, then, organized by the same tree as the pure states. The interpretation of dynamical measurements went even further to suggest that the finite barriers between metastable states were temperature dependent, and apparently diverged into the pure states as the temperature was lowered [142], though of course the approach to divergence could not be accessed experimentally.
@@
Figure 14: Hierarchical organization of metastable states. The coarse-grained free-energy surface is represented at each level corresponding to a given temperature. When temperature is decreased, each valley is subdivided into others. The times $t_{\mathrm{w}}$ and $t_{\mathrm{w}}^{\prime}$ are indicated that are necessary to explore, at $T_{\mathrm{m}}$ and $T_{\mathrm{m}}-\delta T$, respectively, the region of phase space bounded by the same barriers. The closest common ancestor to all states within the space bounded by $\Delta_{1}$ at $T_{\mathrm{m}}$ and $T_{\mathrm{m}}-\delta T$ is the same, and its corresponding value of the overlap function is $q_{1}$. The sketch also shows that, as the system explores more of phase space, it encounters ever-increasing barrier heights, and that the free-energy surface has a self-similar structure. Reproduced from Fig. 6 of Lederman et al. [188].
The dynamics associated with the constructs exhibited in Figs. 13 and 14 follows from their structure. At $t=0$, the temperature is quenched from above $T_{\mathrm{g}}$ in the paramagnetic phase to a measuring temperature $T_{\mathrm{m}}$ in the spin-glass phase. Instantaneously, the spin glass is found in some state $\alpha$ at $T_{\mathrm{m}}$ with self overlap $q_{\alpha\alpha}=q_{\text{\tiny EA}}(T_{\mathrm{m}})$ [Eq. (8)]. As time progresses, the quenched initial state organizes from its paramagnetic structure into a progressively growing spin-glass state, expressed through a growing spin-glass coherence length $\xi_{\mathrm{micro}}(t)$, which is a function of the time elapsed since the quench, termed the waiting time $t_{\mathrm{w}}$. As $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ grows, larger and larger free-energy barrier heights are occupied, having been created by the growing coherence length. The overlap $q_{\alpha\beta}$ with the initial state diminishes and is given by
@@
Experimentally, one needs a finite magnetization to make a measurement. This requires a magnetic field that can polarize the otherwise non-magnetic quenched state. Bouchaud [65] introduced another piece of phenomenology, a Zeeman energy $E_{\text{Z}}$, which reduces each of the barrier heights and is given by [159]
$E_{\text{Z}}=-M_{\text{FC}}H=-\chi_{\text{FC}}N_{\text{s}}H^{2},$ (10)
$N_{\text{s}}=\xi(t_{\mathrm{w}})^{D-\theta/2},$ (11)
where $\theta$ is the replicon exponent [33, see also Eq. 23] and $D$ is the spatial dimension. Thus, combining the reduction of the energetic barrier, Eq. (10), to the peak of the relaxation function $S(t,t_{\mathrm{w}};H)$ of Eq. (2) one obtains:
It is commonplace to age the spin glass over a time $t_{\mathrm{w}}$, then probe its time dependent magnetization at a time $t$ after a magnetic field is turned on at $t_{\mathrm{w}}$, the so-called zero-field-cooled (ZFC) magnetization. Alternatively, the system can be prepared in the presence of a magnetic field, at a time $t$ after which the field is reduced to zero at time $t_{\mathrm{w}}$, the so-called thermoremanent magnetization (TRM); see Sec. I.2. The spin-glass correlation length is then denoted by $\xi(t,t_{\mathrm{w}};H)$.
Experiments can only extract a spin-glass correlation length through Eq. (12), see Sec. II.6. As a consequence, and because of a plethora of different correlation lengths accessible in simulations, we refer to it as $\xi_{\text{Zeeman}}(t,t_{\mathrm{w}};H)$. It will be shown in Sec. V.2 that it is found to be roughly equivalent to the $\xi_{\text{micro}}(t,t_{\mathrm{w}};H)$ that is extracted through simulations in fixed-temperature protocols. The first use of Eqs. (1012) to extract $\xi_{\text{Zeeman}}(t,t_{\mathrm{w}};H)$ can be found in Joh et al. [159]. A subsequent analysis in terms of the phenomenological model discussed here can be found in Paga et al. [231]. The use of Fig. 14 to display both the self similarity of the metastable states between the infinite barriers of the Parisi pure states, and their evolution toward the pure states can be found in Hammann et al. [142].

### v2

@@ I.3 A glossary for aging dynamics
    Two relevant temperatures are $T_{\mathrm{g}}$ and $T_{\mathrm{m}}$. The glass transition occurs at $T_{\mathrm{g}}$, which also is the critical temperature for the phase transition that separates the high-temperature paramagnetic phase from the spin-glass phase that becomes stable below $T_{\mathrm{g}}$. In a typical experimental setting, the system is quickly cooled from some very high temperature $T>T_{\mathrm{g}}$ to the measuring temperature $T_{\mathrm{m}}$.
@@
    Two relevant times are $t$ and $t_{\mathrm{w}}$. We set our initial time as the time for the temperature quench from $T>T_{\mathrm{g}}$ to $T_{\mathrm{m}}<T_{\mathrm{g}}$. The system is allowed to relax at $T_{\mathrm{m}}$ for a waiting time $t_{\mathrm{w}}$. Then, at time $t_{\mathrm{w}}$ the value of the externally applied magnetic field is varied (in the ZFC protocol the field is switched on at time $t_{\mathrm{w}}$, while in the TRM protocol the field is switched off). The magnetization is measured at a later time $t+t_{\mathrm{w}}$, giving us information about the sample’s response to the variation of the magnetic field at the earlier time $t_{\mathrm{w}}$.
@@
    We shall need two kind of averages. The thermal average of a quantity $\mathcal{A}$ will be denoted by $\langle\mathcal{A}\rangle$. The Monte Carlo method [328, 222] is ideally suited to estimate $\langle\mathcal{A}\rangle$. Details specific for nonequilibrium simulations can be found in Belletti et al. [46]; see also Alvarez Baños et al. [15], Martin-Mayor et al. [248] for details about equilibrium simulations and Yllanes [369] for a general discussion. To obtain the average over disorder, which is always carried out after the thermal average, one first independently simulates several systems, each with an independent choice of the random couplings, and then takes the average of the different thermal averages. The average over disorder will be denoted by $\overline{\langle\mathcal{A}\rangle}$. Although the average over disorder is absolutely crucial for numerical simulations, experiments are usually carried out for just one sample. Due to self-averaging, this is usually a sound procedure under experimental conditions (see Sec. II.8).
@@
    The two-time correlation length $\zeta(t,t_{\mathrm{w}})$ describes the growth of order in the aging regime when the two times $t_{\mathrm{w}}$ and $t+t_{\mathrm{w}}$ are compared [186, 22, 46]. This quantity, defined in Sec. IV.2, will be vital for understanding temperature-varying protocols. [^1] Our distinction between coherence and correlation lengths is inspired in the coarsening dynamics of a ferromagnetic system with Ising spins. In this analogy, the coherence length is the size of the magnetic domains while the correlation length indicates the length scale to which statistical correlations propagate inside a domain.
@@
    The experimentally accessible $\xi_{\mathrm{Zeeman}}$ introduced by Joh et al. [188] will be explained in Sec. II.3. This length is related to the lowering of free-energy barriers because of the Zeeman effect and is nonlinear in the magnetic field.
As explained in Sec. V, the relationship between the three length scales $\xi_{\mathrm{micro}}$, $\xi_{\mathrm{Zeeman}}$ and $\zeta$ is the key to the study of the nonlinear response of a spin glass in a magnetic field. A central result will be that $\xi_{\mathrm{micro}}$ and $\xi_{\mathrm{Zeeman}}$ turn out to behave very similarly whenever they can be compared. In other words, it is possible to use bulk measurements to determine apparently microscopic quantities (the size of the spin-glass domains). This correspondence could suggest also referring to $\xi_{\mathrm{Zeeman}}$ as a coherence length, but we shall follow the experimental tradition of referring to $\xi_{\mathrm{Zeeman}}$ as a correlation length.
@@ II.1 The Edwards-Anderson and Sherrington-Kirkpatrick models and overlap as an order parameter
By assuming that $\sum_{i,j}J_{ij}\epsilon_{ij}=0$ on all length scales of the system (where $J_{ij}$ is the strength of the interaction between sites $i$ and $j$ and $\epsilon_{ij}$ is the probability of finding two interacting spins at those positions), Edwards and Anderson [121] introduced a very simplified model of classical spins (which will turn out to embed all relevant features of the problem), the Edwards-Anderson (EA) model. A different approach, based on the presence of potentially magnetized domains, had been proposed before by Adkins and Rivier [2]. The fact that the disordered interactions are distributed at random and cannot change in time —we say that the disorder is quenched— is a crucial element of the model.
@@
where the primed sum runs over the pairs of nearest neighbors in a $D$-dimensional hypercubic lattice. The spins $\sigma_{i}$ can take different forms. For instance, in the Heisenberg spin glass, they are unit vectors in space. Unless otherwise specified, however, in this review, the EA model will always be considered with Ising spins that can only take the values $\pm 1$ (see Sec. VII.1.3 for a discussion of the practical differences between these two models). The applied magnetic field $H$ will be considered to be zero, for now.
Notice that this Hamiltonian is written in units such that $k_{\text{B}}=1$. To compare with experimental results, temperatures can be expressed in terms of $T_{\mathrm{g}}$ (more sophisticated scaling variables will be introduced later). As for the magnetic field, it can be estimated [35, 23], that $H=1$ roughly corresponds to $5\times 10^{4}$ Oe. The renormalization group [366, 377, 296] teaches us that the effect of the magnetic field scales with its value in natural units that depend on the coherence length. This important point will be explored throughout the review. Finally, the natural unit of time in simulations is the lattice sweep, which is equivalent to $\sim 1$ ps [271].
It is crucial that the quenched couplings, which do not change in the course of the dynamics, can have both signs. A possible choice is $J_{ij}=\pm 1$ with a probability of one half, or that they be normally distributed, with zero mean and unitary variance, $P\left(J_{ij}\right)\sim\exp(-{J_{ij}^{2}}/{2J^{2}})$. We will not discuss here the possibility of a nonzero expectation value for the couplings.
Because the quenched couplings can be both positive or negative and are distributed at random, some of the products of the couplings on a closed path of consecutive lattice links will be positive and some will be negative. This implies, in turn, that not all interactions can be satisfied at the same time. The system is frustrated in the sense that all interactions can never be completely satisfied. The energy cannot be as low as it could be in a uniform system. Typically, frustration leads to degeneracy: the low-energy states of such a system have a higher energy than those states found in a pure system, but are more numerous in the disordered system.
In the EA model, because the interactions between spins oscillate in sign as a function of separation, ferromagnetism does not appear at any relevant length scale. On the contrary, for $T$ lower than a critical $T_{\mathrm{g}}$, a spin-glass phase appears, with the spins staying close to fixed directions that appear to be random. This results in overall zero magnetization, but with a nonzero square of the local magnetization leading to the cusp in the susceptibility observed in experiments (see Sec. I.2).
@@
The absence of long-range magnetic order leads to the concept of spin overlap as an order parameter, crucial to all further developments of the field. One considers the spin $\sigma_{i}$ at site $i$ and defines:
$q_{i}\left(t\right)\equiv\lim_{\tau\to\infty}\left\langle\sigma_{i}\left(t\right)\sigma_{i}\left(t+\tau\right)\right\rangle,$ (4)
which can be different from zero in the low-$T$, broken-symmetry phase. The brackets $\langle\cdots\rangle$ signify the average over different thermal histories. A more convenient way of measuring the same overlap is based on the use of two replicas: copies of the system with the same quenched disorder (same random couplings in the Hamiltonian)
@@
This is the most common observable in numerical simulations. The concept of a nonzero overlap of configurations with zero magnetization is illustrated in Fig. 12.
@@
As with the overlap, we can associate a $P(q_{\text{l}})$ to this observable. The behavior of the overlap and the link overlap under spin flip is very different: consider two configurations of spins with the same disorder which differ in a domain of size $L$, where the spins have flipped. The link overlap provides an estimate for the surface of the domain, whereas the overlap reflects its volume.
The EA spin glass, an Ising-like finite-dimensional model with quenched, frustrated interactions, cannot be solved exactly and cannot be dissected too deeply by analytical techniques. The experience of the Ising ferromagnet suggested introducing the mean-field, nonlocal Sherrington-Kirkpatrick (SK) model [326, 212].
@@
This mean-field model has the usual paramagnetic phase at high temperatures. The delicate point, which will be the source of trouble and delight, is that, because the couplings are quenched, we have to compute the disorder average not of the partition function $Z$ (which would anneal the disorder degrees of freedom) but of its logarithm, $\ln Z$. That is, we average physical observables in a given disorder realization (a sample) and then we average these thermal expectation values over the disorder. To do this, EA and SK use the identity (the so-called replica trick)
@@
Kirkpatrick and Sherrington [212] refined their analysis two years after their first, groundbreaking paper [326]. They described the same situation, but with an analysis based on the assumption of replica symmetry (RS). Not all was well, however. The solution of the RS ansatz not only disagrees with Monte Carlo simulations but also leads to a nonsensical result, namely a negative value at low $T$ for the entropy, which is a semi-definite positive quantity (proving the inadequacy of the replica-symmetric ansatz). Before this second SK paper, Thouless et al. [343] had introduced a different approach, avoiding the replica trick, that shed further light on the low-$T$ behavior of the SK model. Their so-called TAP equations, together with their many variations, have been for many years and continue to be important for the analysis of spin-glass physics. When the temperature decreases below a critical value $T_{\mathrm{g}}$, the correct solution found at high $T$ under the assumption of symmetry between replicas turns out to be undoubtedly wrong.
@@
The above-described mean-field approximation is only correct above the so-called upper critical dimension $D_{\text{u}}$, which for spin glasses is $D_{\text{u}}=6$ [78]. On the other hand, below the lower critical dimension $D_{\text{l}}$, the low-temperature phase disappears. It is believed that $D_{\text{l}}=5/2$ and it is clear that $2<D_{\text{l}}<3$ [146, 236]; for our purposes what matters is that $D=3$ lies in the nontrivial region. There has been an intense quest for a successful extension of the Parisi theory to $D=3$ in the renormalization-group style [see for example 113, 110]. Because of the infinity of states that appear in Parisi’s solution for the mean-field theory, renormalizing the system is a daunting task. Today, after close to 50 years of studies, there still remains much need for clarification. The interested reader can start from De Dominicis et al. [113] and De Dominicis and Giardina [110], then continue with the following papers and included references: Bray and Moore [73, 74], Bray and Roberts [78], De Dominicis and Kondor [111, 112], Temesvári et al. [338], De Dominicis et al. [109], Temesvári [337].
@@ II.2 Main theoretical pictures for the low-temperature phase
  • Trivial-Non-Trivial (TNT). In the droplet model (in the absence of a magnetic field) the probability distribution $P(q)$ of the overlap $q$ between two pure states, and the link overlap, are both trivial (i.e., $P(q)$ is composed of two Dirac delta functions and one Dirac delta function, respectively). In the TNT approach, $P(q)$ exhibits a nontrivial probability distribution, whereas the link overlap exhibits a trivial probability distribution [217, 291].
  • Replica-Symmetry Breaking (RSB). In this scenario, which consists in extending the features of the mean-field solution to finite dimensions [242, 110], the spin glass has infinite pure states not related by any kind of symmetry. However, this infinite set of states can be organized in an ultrametric fashion. Furthermore, the excitations of the ground state are space-filling. Finally, in this theoretical approach, the spin-glass phase is stable to small magnetic fields (i.e., there exists a phase transition between a paramagnetic phase and a spin-glass phase in the presence of a magnetic field).
Figure 13: The main quantitative difference between the competing theoretical pictures of spin-glass dynamics is related to the value of an exponent $\theta$, the replicon, see Eq. (25) and Marinari et al. [242, 241]. This exponent is nonzero for RSB and zero in the droplet picture, at least in the long-time limit. When represented as a function of a scaling variable $x$ [a properly rescaled $1/\xi(t_{\mathrm{w}},T)$, see 36], the values of $\theta$ obtained for several temperatures fall on a single curve. In addition we show a droplet extrapolation [dashed line, $\theta(x\!\to\!0)\!\propto\!x^{\iota}$, with $\iota\!\approx\!0.15$] and an RSB extrapolation [full line, $\theta(x\!\to\!0)\!>\!0$]. Even the droplet extrapolation, which has $\theta(x)\to 0$ as $x\to 0$, predicts sizable values of $\theta$ for the values of $x$ relevant to experimental length scales (black dots). Data taken from Baity-Jesi et al. [36].
@@ II.3 Phenomenological dynamical model on an ultrametric tree
There are no dynamics associated with pure states in that the free-energy barriers between them are infinite. Experimentally though, time-dependent phenomena are observed. It takes a large amount of chutzpah to suggest that between the pure states there could exist states between which the free-energy barriers were finite and even more to suggest that these states between infinite barriers were self-similar. No matter the temperature, the same ultrametricity holds for metastable states, which are organized by the same tree as the pure states. The interpretation of dynamical measurements goes even further to suggest that the finite barriers between metastable states are temperature dependent, and apparently diverge into the pure states as the temperature is lowered [166], though of course the divergence can not be accessed experimentally.
@@
Figure 15: Hierarchical organization of metastable states. The coarse-grained free-energy surface is represented at each level corresponding to a given temperature. When temperature is decreased, each valley is subdivided into others. The times $t_{\mathrm{w}}$ and $t_{\mathrm{w}}^{\prime}$ are indicated that are necessary to explore, at $T_{\mathrm{m}}$ and $T_{\mathrm{m}}-\delta T$, respectively, the region of phase space bounded by the same barriers. The closest common ancestor to all states within the space bounded by $\Delta(T_{\text{m}})$ and $\Delta(T_{\text{m}}-\delta T)$ is the same, and its corresponding value of the overlap functions are $q_{\text{EA}}(T_{\text{m}})$ and $q_{\text{EA}}(T_{\text{m}}-\delta T)$. The sketch also shows that, as the system explores more of phase space, it encounters ever-increasing barrier heights, and that the free-energy surface has a self-similar structure. Reproduced from Fig. 6 of Lederman et al. [224].
The dynamics associated with the constructs exhibited in Figs. 14 and 15 follow from their structure. At $t=0$, the temperature is quenched from above $T_{\mathrm{g}}$ in the paramagnetic phase to a measuring temperature $T_{\mathrm{m}}$ in the spin-glass phase. Instantaneously, the spin glass is found in some state $\alpha$ at $T_{\mathrm{m}}$ with self overlap $q_{\alpha\alpha}=q_{\text{\tiny EA}}(T_{\mathrm{m}})$ [Eq. (8)]. As time progresses, the quenched initial state organizes from its paramagnetic structure into a progressively growing spin-glass state, expressed through a growing spin-glass coherence length $\xi_{\mathrm{micro}}(t)$, which is a function of the time elapsed since the quench, termed the waiting time $t_{\mathrm{w}}$. As $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ grows, larger and larger free-energy barrier heights are occupied, having been created by the growing coherence length. The overlap $q_{\alpha\beta}$ with the initial state diminishes and is given by
@@
The main consequences of this framework will be considered in Sec. II.6, where the extraction of $\xi$ from experiment is described. An analysis in terms of the phenomenological model discussed here can be found in Paga et al. [288]. The use of Fig. 15 to display both the self similarity of the metastable states between the infinite barriers of the Parisi pure states, and their evolution toward the pure states as the temperature is lowered, can be found in Hammann et al. [166].